An actual yield of 18g, the theoretical yield is unknown. Calculate the theoretical yield?

1
Expert's answer
2022-01-28T06:36:01-0500

Question is incomplete but,


Actual yield is normally smaller than the theoretical yield

Consider burning 35g of glucose


"C_6H_{12}O_6+6O_2\to6CO_2+6H_2O"


Theoretical yield of water can be calculated as follows:

Moles of "C_6H_{12}O_6=\frac{35}{180}"


Moles of water "=6×\frac{35}{180}"


Mass of water "=6×\frac{35}{180}×18"

"=21g"



If 18g of water was produced, instead of 21g


21g represent the theoretical yield.